OpenStudy (anonymous):

write the formula for g(x) as a function of x. f(x)=x^3+5x+1, g(x)=5+f(2x)

OpenStudy (anonymous):

wherever you have x on g(x) plug in f(x)

OpenStudy (anonymous):

\[g(x)=5+f(2x)=5+(2x)^3+5(2x)+1\] which you can clean up with some algebra